instantaneous velocity for a displacement function d(t)= 2-2t at any time is given by


1
Expert's answer
2021-09-12T19:14:15-0400

Gives

Displacement function d(t)=2-2t

Assume d(t)=x=2-2t

Now velocity

"v=\frac{dx}{dt}"

Put value


"v=\frac{d}{dt}(2-2t)=0-2=-2m/sec"

Velocity

"v=-2m\/sec"